OpenStudy (anonymous):

Solve for x. -3 + 8x - 5 = -8

OpenStudy (diyadiya):

-3 + 8x -5 =-8 it can be written as -3 -5 +8x = -8 can you do -3-5 ?

OpenStudy (diyadiya):

@DerrickRose -3-5= ?

OpenStudy (anonymous):

-8

OpenStudy (diyadiya):

Right so now the equation is ?

OpenStudy (anonymous):

A. -2 B. -1 C. 3 - 4 D. 0

OpenStudy (diyadiya):

-3 -5 +8x = -8 -8 + 8x =-8 ok ?

OpenStudy (diyadiya):

Now add 8 on both sides what do you get ?

OpenStudy (anonymous):

-16?

OpenStudy (anonymous):

0??????????????

OpenStudy (diyadiya):

-8 + 8x + 8 =-8+8 Yes its 0 so how will be the equation now ?

OpenStudy (anonymous):

idk o.o

OpenStudy (anonymous):

D

OpenStudy (diyadiya):

@Ala123 Please read the code-of-conduct its against the rules of OS to give away the final answer without any explaination. http://openstudy.com/code-of-conduct @derrickRose -8 + 8x + 8 = -8 + 8 8x = 0 therefore x=0 Hence the answer is D

OpenStudy (anonymous):

Yes sir!